1.Who knows the game sipa? A Physical Education major sought to determine whether all Filipinos are familiar with this game. Among 1,500 respondents, 358 knew the game. Use 95% confidence to estimate the population proportion p and q. Interpret the sample data.

A. Find (za/2)2 given each of the following:

-----------

E


a. With 90% confidence, E = 0.01.

b. With 90% confidence, E = 0.02.

c. With 95% confidence, E = 0.15.

d. With 95% confidence, E = 0.08

e. With 99% confidence, E = 0.05.





1
Expert's answer
2021-04-13T12:23:00-0400

1.

"E=Z_{\alpha/2}\sqrt{\frac{p(1-p)}{n}}"

"n=1500,p=358/1500=0.2387=23.87\%,\ q=76.13\%"

"E=1.96\sqrt{\frac{0.2387\cdot0.7613}{1500}}=0.0216=2.16\%"

The population proportion p and q are in 95% confidence interval with error 2.16%


A. a)

"(Z_{\alpha/2})^2=(1.65\pm0.01)^2"

"2.69\leq (Z_{\alpha/2})^2\leq 2.76"


b)

"(Z_{\alpha/2})^2=(1.65\pm0.02)^2"

"2.66\leq (Z_{\alpha/2})^2\leq 2.79"


c)

"(Z_{\alpha/2})^2=(1.96\pm0.15)^2"

"3.28\leq (Z_{\alpha/2})^2\leq 4.45"


d)

"(Z_{\alpha/2})^2=(1.96\pm0.08)^2"

"3.53\leq (Z_{\alpha/2})^2\leq 4.16"


e)

"(Z_{\alpha/2})^2=(2.58\pm0.05)^2"

"6.40\leq (Z_{\alpha/2})^2\leq 6.92"